# NuRF: Nudging the Particle Filter in Radiance Fields for Robot Visual Localization

### v1

@@ Abstract
Can we localize a robot in radiance fields only using monocular vision? This study presents NuRF, a nudged particle filter framework for 6-DoF robot visual localization in radiance fields. NuRF sets anchors in SE(3) to leverage visual place recognition, which provides image comparisons to guide the sampling process. This guidance could improve the convergence and robustness of particle filters for robot localization. Additionally, an adaptive scheme is designed to enhance the performance of NuRF, thus enabling both global visual localization and local pose tracking. Real-world experiments are conducted with comprehensive tests to demonstrate the effectiveness of NuRF. The results showcase the advantages of NuRF in terms of accuracy and efficiency, including comparisons with alternative approaches. Furthermore, we report our findings for future studies and advancements in robot navigation in radiance fields.
@@ I Introduction
Visual localization on a map is a critical topic for robot navigation, which is close to how humans perceive the world. Existing feature-based techniques, like Perspective-n-Point (PnP) [1] offer stable and accurate pose tracking for robots, they struggle to directly address global localization in cases of localization failure [2]. On the other hand, learning-based approaches like visual place recognition (VPR) [3, 4, 5] can regress or retrieve poses through trained networks from scratch. These approaches fall short of providing a comprehensive and cost-effective solution for continuous 6-degree-of-freedom (6-DoF) pose tracking.
There is a growing need for a unified framework that integrates global localization and pose tracking for robot visual localization, enabling the use of one localization method on one map for two tasks. Particle filters, also referred to as Monte Carlo localization [6], are commonly employed as back-end estimators for such problems [7] due to their ability to represent arbitrary distributions of poses. However, achieving effective 6-DOF pose estimation localization often requires a large number of particles, leading to inefficiencies in visual localization for robots. Furthermore, explicit visual features or space discretization are typically necessary to support particle filter-based localization, which restricts the practicality of real-world deployment.
@@
In recent years, the neural radiance field (NeRF) has emerged as a novel implicit representation [8], enabling the generation of high-quality, dense photographic images without the need for intricate feature extraction and alignment. This has led to the exploration of using NeRF-based map representations in pose estimation tasks. Maggio et al. [9] proposed to use particle filter as a back-end estimator for pose estimation from radiance field-based maps. Our previous work by Lin et al. [7] utilized similar techniques, incorporating VPR for pose estimation. However, current methods primarily rely on pixel-level differences for estimating robot positions, disregarding the valuable semantic and feature information present in the rendered images. This limitation results in slow convergence, particularly in global localization tasks.
To address these limitations, we design a nudged particle filter framework in radiance fields (NuRF), which achieves robot visual localization only using NeRF as map. NuRF incorporates the nudged technique and adaptive image resolution, thus enhancing global localization and pose tracking performances. Specifically, it generates a diverse set of 2D submanifold images on the special Euclidean group (SE(3)) and vectorizes them using a place recognition network. This process creates a comprehensive vector database that encapsulates the rich information present in the radiance fields. This vector database forms the foundation of our particle filter, representing the posterior probability density of the robot’s pose through particle distributions and weights. By utilizing this database, NuRF is able to estimate and track the robot pose more robustly and accurately.
In summary, our primary contributions are threefold:
  • We set anchor poses for visual place recognition, improving the convergence of particle filters by nudging.
  • We design an adaptive scheme that switches between global localization and position tracking, thus achieving robust and efficient indoor visual localization.
@@
The rest of our paper is organized as follows: Section II introduces the related work on radiance field, visual place recognition, and nudged particle filter; Section III presents the problem formulation and system overview. We detail the proposed method in Section IV and demonstrate its effectiveness with real-world experiments in Section V. The conclusions and future studies are summarized in Section VI.
@@ II Related Work
II-A Radiance Field
A radiance field is a representation of how light is distributed in three-dimensional space, capturing the interactions between light and surfaces in the environment [10]. In the context of neural radiance fields, i.e., NeRF, it allows us to generate images from a neural model by giving any given camera pose. Hence, the neural radiance field can be seen as a low-dimensional sub-manifold that encompasses all the rendered images of the scene in a high-dimensional pixel space. In recent years, NeRF is a new wave for robotic applications, as summarized in the survey paper [11].
Loc-NeRF by Maggio et al. [9] introduced a Monte Carlo localization method for estimating the posterior probability of individual camera poses in a given space by rendering multiple random images within a radiance field. These rendered images are then compared to the target image in pixel space, evaluating the pixel-space distances. The recent advancement in explicit neural radiance field Gaussian splatting [12] has achieved a rendering rate of 160 frames per second (fps) per image while requiring less than 500MB of storage [13]. Moreover, this technology can be implemented using handheld devices such as smartphones [14]. In the 3DGS-ReLoc by Jiang et al. [15], Gaussian splatting representation was utilized for visual re-localization in urban scenarios, following a coarse-to-fine manner for global localization. Specifically, the camera was first located by similarity comparisons, and then refined by the PnP technique.
This study is inspired by the Loc-NeRF [9]. Their work demonstrated that robot localization can be performed with a radiance field-based map. However, it is worth noting that the evaluation of the algorithm’s localization performance was limited to a confined space. Even using simulated annealing for acceleration, it still requires over 400 hours to achieve a position estimation error of 0.5 $m$ within a 4 $m^{2}$ room. This indicates that the proposed method may not be practically feasible for engineering implementation.
@@ II-B Visual Place Recognition
Visual place recognition (VPR) is a fundamental problem in robotics and computer vision [3], aiming to provide a coarse pose estimation by retrieving geo-referenced frames in close proximity. Modern learning-based VPR solutions, such as AnyLoc [16], offer comprehensive and accurate frame retrieval due to their robust representation of both images and maps. However, VPR achieves localization by considering the reference pose approximating the query pose. This inherent approximation limits the localization accuracy, especially when there is a significant deviation between the current pose and the queried poses in the database. Consequently, VPR is typically employed as a coarse or initial step in high-precision localization systems [17], which is also analyzed in the survey paper for global LiDAR localization [2]. On the other hand, constructing a database of landmark anchors for reference purposes can be a time-consuming and labor-intensive process. It involves collecting a reference image database from the robot’s onboard camera and odometer data as it navigates through the environment.
@@ II-C Nudged Particle Filter
Nudging refers to a sampling approach in particle filtering that guides particles towards regions of high expected likelihood [18]. This technique is particularly effective when the posterior probabilities concentrate in relatively small areas of the state space, making it well-suited for high-resolution observation models, such as images. In our previous work conducted by Lin et al. [7], we conducted experiments demonstrating that the utilization of nudging particles enabled accurate tracking of a robot’s pose, even in cases where the initial distribution was incorrect and had low variance. However, we did not achieve successful global localization using the nudged particle filter in that particular study.
III Problem Formulation and System Overview
Mathematically, both implicit and explicit radiance fields can be represented as a pair, denoted as $[\mathcal{G},L]$. The term $\mathcal{G}$ represents a 3D scene in the world frame, and $L$ is the rendering function that takes $\mathcal{G}$ and a pair of camera parameters $[K,T]$ as inputs and produces the rendered image $I_{\mathcal{G}}$. The rendering function $L$ can be formulated as:
$I_{\mathcal{G}}=L(\mathcal{G},K,T)$ (1)
in which camera parameter $K$ is the intrinsic matrix that encodes the field of view and the resolution of the rendered image, and the extrinsic matrix $T\in$ SE(3) represents the 6-DoF pose of the camera regarding the world frame. For convenience, let the pose of a camera-mounted robot at time $\tau$ be the same as the extrinsic matrix of the camera $T_{\tau}$. From time $\tau-1$ to $\tau$, the robot pose is integrated by relative transformations $H_{\tau}\in$ SE(3), described as follows:
@@ III Problem Formulation and System Overview
in which $\boldsymbol{t},\boldsymbol{u}\in\mathbb{R}^{3}$ are the translation vectors and $\boldsymbol{R},\boldsymbol{V}\in\mathbb{R}^{3\times 3}$ are rotation matrices. For two consecutive poses, we assume the sensors equipped on the robot can capture two images $I_{\tau-1}$, $I_{\tau}$ and the motion $\hat{H}_{\tau}$. Note that the motion model is perturbed by noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\Sigma)\in\mathbb{R}^{3}$ and $\boldsymbol{\varepsilon}\sim\boldsymbol{U}[-a,a]\in\mathbb{R}^{3}$. The relative transformation is therefore given by:
$\hat{\boldsymbol{u}}_{\tau}=\boldsymbol{u}_{\tau}+\boldsymbol{\epsilon}$ (3)
$\hat{\boldsymbol{V}}_{\tau}=f_{\boldsymbol{R}}(\boldsymbol{\varepsilon})\boldsymbol{V}_{\tau}$
where $f_{\boldsymbol{R}}$ converts the sampled Euler angle to a rotation matrix. Then, the posterior probability density function of the robot pose $T_{\tau}$ can be written as:
$p(T_{\tau}|I_{1:\tau},\hat{H}_{2:\tau})=\frac{p(I_{1:\tau}|\hat{H}_{2:\tau},T_{\tau})p(\hat{H}_{2:\tau},T_{\tau})}{p(I_{1:\tau},\hat{H}_{2:\tau})}$ (4)
Then, we can formulate the visual-based global localization problem as Problem III.1.
@@
Given a sequence of RGB images $\boldsymbol{I}=[I_{1},I_{2},\dots,I_{\tau}]$ and a motion model $\boldsymbol{\hat{H}}=[\hat{H}2,\hat{H}3,\dots,\hat{H}{\tau}]$, the task is to approximate the posterior distribution of the robot pose $T{\tau}$ as delineated in equation (4), and to achieve the maximum likelihood estimation of the robot pose, denoted as $T^{*}_{\tau}$.
(a) X direction
(b) Y direction
(c) Z direction
Figure 2: We assess the similarity between a query image and images rendered in the same orientation but shifted along the X/Y/Z axis using pixel-wise similarity (SSIM) and feature-wise similarity (LPIPS).
Let $\varXi_{\tau}=\{\zeta_{\tau,i}\}^{N}_{i=1}$ be the set of particles at time $\tau$, where each particle $\zeta_{\tau,i}=[w_{\tau,i},T_{\tau,i}]$ with weight $w_{\tau,i}$ and pose $T_{\tau,i}$. Consequently, the posterior probability density function (4) can be approximated as follows:
@@
in which $\delta$ is the Dirac delta function, and $\eta$ is a normalisation constant; if $T_{\tau,i}$ is a sample point in (4) and $w_{\tau,i}$ presents the likelihood that the camera captured image $I_{\tau}$ at $T_{\tau,i}$ after motions $\hat{H}_{2:\tau}$ as:
$w_{\tau,i}=p(I_{\tau}|\hat{H}_{2:\tau},T_{\tau})=\frac{p(I_{1:\tau}|\hat{H}_{2:\tau},T_{\tau})}{p(I_{1:\tau-1}|\hat{H}_{2:\tau},T_{\tau})}$ (6)
It’s very difficult to calculate (6) directly. But, with the radiance field representation, we could approximate (6) using the similarity between the rendered images and the observed image at $T_{\tau,i}$. The most intuitive method to calculate the similarity between two images is calculating the pixel-wised distances between those images. The structural similarity index measure (SSIM) [19] is a typical pixel-wised image similarity function. With SSIM, the distance between image $x$ and $y$ can be calculate as:
$\text{Distance}(x,y)$ $=1-\text{SSIM}(x,y)$ (7)
$=1-\frac{(2\mu_{x}\mu_{y}+c_{1})(2\sigma_{xy}+c_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+c_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2})}$
in which $\mu$ and $\sigma$ represent the mean and standard deviation of each image, respectively. Another kind of method involves learning techniques that transform the image into a learned feature space and subsequently utilize the distances within this feature space to estimate the similarity. A typical approach is learned perceptual image patch similarity (LPIPS)[20]:
$\text{Distance}(x,y)$ $=\text{LPIPS}(x,y)$ (8)
$=\sum_{l}^{L}\frac{\sum_{h,w}||w_{l}\odot(\hat{x}^{l}_{hw}-\hat{y}^{l}_{h}w)||^{2}_{2}}{H_{l}W_{l}}$
in which $\hat{x}^{l}$ and $\hat{y}^{l}$ represent the feature maps of the images at a particular feature layer, and $w_{l}$ denotes the corresponding layer weight.
@@
Figure 2 shows that the linear correlation between image similarity and robot pose is more prominent when utilizing the feature-wised approach [21]. This attribute makes it easier to employ the feature-wised similarity as the key factor to approximate the term in (6). However, although it is more accurate to use the feature-wise similarity function to weight particles, projecting the image rendered by each particle into the feature space needs large computing resources. Therefore, a pixel-based similarity function is required if a real-time particle filter is needed. This leads to Problem 9.
Problem III.2
Given two particle sets $\varXi^{1}$ and $\varXi^{2}$. Known that $T^{1}\equiv T^{2}$, and $w^{1}$ and $w^{2}$ are calculated by (7) and (8) respectively. Add some extra particles into $\varXi^{1}$ to get $\varXi^{1+}$. And make $\varXi^{1+}$ satisfied:
$\varXi^{1+}=\arg\min_{\varXi^{1+}}\parallel 1-\frac{f(z;\varXi^{2})}{f(z;\varXi^{1+})}\parallel$ (9)
The solution to Problem 9 holds the key to effectively addressing Problem III.1 for practical deployment. Furthermore, in practice, another classic problem in robot localization is the kidnapping problem, which can be described as Problem III.3.
Problem III.3
Move the robot to a new location and observe a new image $I_{\tau+1}$, but incorrect motion information is provided. Detect this motion in limited time steps.
To address the outlined problems, we introduce the NuRF framework as depicted in Figure 3. This framework integrates anchor points in close proximity to the target image $I_{\tau}$ within the feature space during the resampling phase. By implementing this strategy, both Problem 9 and Problem III.1 are effectively addressed for real-world deployment. Additionally, the utilization of nudging particles facilitates the monitoring of particle dispersion, i.e., the variance of particles, which is used to switch the global localization and poser tracking in the radiance fields. We will detail the NuRF modules in the following section.

### v2

@@ Abstract
Can we localize a robot on a map only using monocular vision? This study presents NuRF, an adaptive and nudged particle filter framework in radiance fields for 6-DoF robot visual localization. NuRF leverages recent advancements in radiance fields and visual place recognition. Conventional visual place recognition meets the challenges of data sparsity and artifact-induced inaccuracies. By utilizing radiance field-generated novel views, NuRF enhances visual localization performance and combines coarse global localization with the fine-grained pose tracking of a particle filter, ensuring continuous and precise localization. Experimentally, our method converges 7 times faster than existing Monte Carlo-based methods and achieves localization accuracy within 1 meter, offering an efficient and resilient solution for indoor visual localization.
@@ I Introduction
Visual localization on a map is a critical topic for robot navigation, which is close to how humans perceive the world. An ideal visual localization system is desired to accurately localize robots not only globally but also continuously [1]. Existing feature-based techniques, like perspective-n-point (PnP) [2] offer stable and accurate continuous pose tracking for robots, they struggle to directly address global localization in cases of localization failure [3]. On the other hand, learning-based approaches like visual place recognition (VPR) [4, 5, 6] can regress or retrieve poses through trained networks from scratch. These approaches fall short of providing a comprehensive and cost-effective solution for continuous 6-degree-of-freedom (6-DoF) pose tracking. There is a growing need for a unified framework that integrates global localization and pose tracking for robot visual localization, enabling the use of one localization method on one map for two tasks.
Recent advancements in radiance field rendering, particularly neural radiance fields (NeRF)[7] and 3D Gaussian splatting (3DGS)[8], have shown significant potential in overcoming the limitations of traditional visual localization. Radiance field-based maps can enhance visual-based localization by providing dense, high-quality, and renderable 3D reconstructions of environments [9, 10, 11, 12], thereby improving localization capabilities in texture-less or occluded regions. Recent studies [10, 11] demonstrate the feasibility of utilizing radiance fields for visual global localization. However, casually captured radiance fields often suffer from unreal artifacts, such as floaters, which can emerge and are not part of the real view [13]. These ghostly artifacts critically impact visual localization in terms of robustness and accuracy. Existing regularizers [14, 15, 16] are ineffective in detecting or eliminating floaters; alternative optimization methods based on 3D diffusion models [13] are too time-consuming for real-time robot localization.
@@
To address the limitations above, we propose an adaptive nudged particle filter in radiance field (NuRF). NuRF enhances VPR through radiance field representations and nudges VPR information to mitigate the impact of floaters. First, we utilize the radiance field to generate novel views from new perspectives, overcoming the challenge of sparseness for VPR. Second, inspired by a previous work [17], the nudged particle filter fuses the coarse global localization information from VPR with the particle filter’s fine-grained pose estimation, enabling continuous, global and efficient localization. Third, this hybrid approach also ensures a smooth transition from global localization to accurate pose tracking via an adaptive mechanism. This approach effectively balances computational efficiency and precision, addressing the challenges of robot localization using NeRF as maps. Overall, the contributions are summarized as follows:
  • We achieve a monocular robot localization by using radiance fields as maps and leveraging VPR techniques.
  • We propose nudging the particle filter to enhance visual localization performance and overcome the challenge of artifacts in radiance fields.
  • We design an adaptive scheme that switches between global localization and position tracking for robust and efficient visual localization.
@@
The rest of our paper is organized as follows: Section II introduces the related work on radiance field, visual place recognition, and nudged particle filter; Section III presents the problem formulation. We detail the proposed method in Section IV and demonstrate its effectiveness with real-world experiments in Section V. The conclusions and future studies are summarized in Section VI.
@@ II Related Work
II-A Localization in Radiance Fields
A radiance field is a representation of how light is distributed in three-dimensional space, capturing the interactions between light and surfaces in the environment [7]. It can be seen as a low-dimensional sub-manifold that encompasses all the rendered images of the scene in a high-dimensional pixel space. In recent years, radiance field is a new wave for robotic applications, as summarized in the survey paper [18].
Loc-NeRF by Maggio et al. [10] introduced a Monte Carlo localization method for estimating the posterior probability of individual camera poses in a given space by rendering multiple random images within a radiance field. These rendered images are then compared to the target image in pixel space, evaluating the pixel-space distances. The recent advancement in explicit neural radiance field Gaussian splatting [8] has achieved a rendering rate of 160 frames per second (fps) per image while requiring less than 500MB of storage [19]. Moreover, this technology can be implemented using handheld devices such as smartphones [20]. In the 3DGS-ReLoc by Jiang et al. [12], Gaussian Splatting representation was utilized for visual re-localization in urban scenarios, following a coarse-to-fine manner for global localization. Specifically, the camera was first located by similarity comparisons, and then refined by the PnP technique. This indicates that the proposed method might not be practically feasible for engineering implementation.
@@ II-B Visual Place Recognition
VPR is a fundamental problem in robotics and computer vision [4], aiming to provide a coarse pose estimation by retrieving geo-referenced frames in close proximity. Modern learning-based VPR solutions, such as AnyLoc [21], offer comprehensive and accurate frame retrieval due to their robust representation of both images and maps.
However, VPR achieves localization by considering the reference pose approximating the query pose. This inherent approximation limits the localization accuracy, especially when there is a significant deviation between the current pose and the queried poses in the database. Consequently, VPR is typically employed as a coarse or initial step in high-precision localization systems [22], which is also analyzed in the survey paper for global LiDAR localization [3]. On the other hand, constructing a database of landmark anchors for reference purposes can be a time-consuming and labor-intensive process. It involves collecting a reference image database from the robot’s onboard camera and odometer data as it navigates through the environment.
@@ II-C Nudged Particle Filter
Nudging refers to a sampling approach in particle filtering that guides particles towards regions of high expected likelihood [23]. This technique has been demonstrated to effectively address issues with Bootstrap particle filters [24] when modeling errors are present, and it’s particularly effective when the posterior probabilities concentrate in relatively small areas of the state space, making it well-suited for high-resolution observation models, such as images. In the previous work conducted by Lin et al. [17], conducted experiments demonstrating that the utilization of nudging particles enabled accurate tracking of a robot’s pose, even in cases where the initial distribution was incorrect and had low variance.
The preliminary version of Nudged Particle Filter was developed with the objective of rectifying erroneous kinetic modeling assumptions through the use of nudging and did not achieve successful global localization using the nudged particle filter in that particular study. We initially proposed the application of Nudged Particle Filter to address deficiencies in observation models like Radiance Field and global localization.
III Problem Formulation
The visual localization problem can be formulated as an estimation of an unknown posterior distribution. For convenience, we will use the special Euclidean group (SE(3)) to present the camera pose or camera extrinsic, let the pose of a camera-mounted robot at time $\tau$ be the same as the extrinsic matrix of the camera $T_{\tau}\in$ SE(3). From time $\tau-1$ to $\tau$, the robot pose is integrated by relative transformations $H_{\tau}\in$ SE(3), described as follows:
@@ III Problem Formulation
in which $\boldsymbol{t},\boldsymbol{u}\in\mathbb{R}^{3}$ are the translation vectors and $\boldsymbol{R},\boldsymbol{V}\in\mathbb{R}^{3\times 3}$ are rotation matrices. For two consecutive poses, we assume the sensors equipped on the robot can capture two images $I_{\tau-1}$, $I_{\tau}$ and the motion $\hat{H}_{\tau}$. Mathematically, both implicit and explicit radiance fields model can be represented as a function $\{L:\text{SE(3)}\mapsto\mathbb{R}^{W\times H}\}$. It provides a mapping of the extrinsic matrix $T_{\tau}$ to a $W\times H$ image $\hat{I}_{\tau}$ that with minimum render loss $\text{LOSS}(\cdot,\cdot)$.
$L(T_{\tau})=\arg\min_{\hat{I}_{\tau}}\text{LOSS}(I_{\tau},\hat{I}_{\tau})$ (2)
Then, the posterior probability density function of the robot pose $T_{\tau}$ can be written as:
$p(T_{\tau}|I_{1:\tau},\hat{H}_{2:\tau},L)=\frac{p(I_{1:\tau}|L,T_{\tau})p(T_{\tau}|\hat{H}_{2:\tau})}{p(I_{1:\tau}|L)}$ (3)
and the visual-based localization problem can be described as an estimation problem for unknown distribution, described as follows:
@@
Given a sequence of RGB images $\boldsymbol{I}=[I_{1},I_{2},\dots,I_{\tau}]$, motion sequence $\boldsymbol{\hat{H}}=[\hat{H}_{2},\hat{H}_{3},\dots,\hat{H}_{\tau}]$, and the radiance field model $L$, how to approximate the posterior distribution of the robot pose $T{\tau}$ as described in Equation (3).
Problem III.1 is a widely studied problem in the robotics literature for unknown distribution estimation and has been solved by Sequential Monte Carlo (SMC) commonly [25]. Let $\varXi_{\tau}=\{\zeta_{\tau,i}\}^{N}_{i=1}$ be the set of particles at time $\tau$, where each particle $\zeta_{\tau,i}=[w_{\tau,i},T_{\tau,i}]$ with weight $w_{\tau,i}$ and pose $T_{\tau,i}$. Consequently, the posterior probability density function (3) can be approximated as follows:
@@
in which $\delta$ is the Dirac delta function, and $\eta$ is a normalization constant. The state transition probability $p(T_{\tau}|\hat{H}_{2:\tau})$ is given by Equation (1), to minimize the diversity between $f(z;\varXi_{\tau})$ and $p(T_{\tau}|I_{1:\tau},\hat{H}_{2:\tau},L)$, the un-normalized importance weights based on the likelihood should satisfied [26]:
$w_{\tau,i}=w_{\tau-1,i}p(I_{\tau}|L,T_{\tau,i})$ (5)
and the likelihood $p(I_{\tau}|L,T_{\tau,i})$ can be computed directly by the pixel-wised render loss function $\text{LOSS}(\cdot,\cdot)$ in Equation (2).
As presented in Equation (2), the radiance field rendering function $L$ is optimized via the pixel-wise loss function $\text{LOSS}(\cdot,\cdot)$, indicating that the weights derived from the loss function could be affected by these artifacts. In this paper, we consider it is possible to eliminate unrealistic artifacts by representing images using advanced visual features (descriptors) [21], thereby retaining only the visual descriptors of the environment. Consequently, a further key focus of this study is how to selectively adjust a small subset of particles to decrease the feature loss based on VPR information during the particle filter update process to enhance estimation accuracy. This consideration leads to the following Problem 6, where the VPR loss is defined by the cosine distance between the visual feature vector of the rendered image and the observed image.
Problem III.2
Given a particle set $\varXi$ and $w$ is calculated by pixel-wised loss. The enhancement of the particle set to better approximate the $p(T_{\tau}|I_{1:\tau},\hat{H}_{2:\tau},L)$ involves minimizing the discrepancy between true distribution and the particle-based estimate $f(z;\varXi^{+})$:
$\varXi^{+}=\arg\min_{\varXi^{+}}\parallel 1-\frac{p(T_{\tau}|I_{1:\tau},\hat{H}_{2:\tau},L)}{f(z;\varXi^{+})}\parallel$ (6)